# Co-Evolving Agents: Learning from Failures as Hard Negative

### v1

@@ 4.3 Co-evolutionary Training
The target agent improves by incorporating failure trajectories generated by the failure agent into its preference optimization. We construct a preference dataset $\mathcal{D}_{\text{tgt}}$ consisting of three types of trajectory pairs: (i) expert trajectories versus target-predicted trajectories, (ii) expert trajectories versus failure-agent trajectories, and (iii) failure trajectories from the target versus those from the failure agent. Formally,
$\mathcal{D}_{\text{tgt}}=\Big\{(u,e_{\text{chosen}},e_{\text{rejected}})\,\Big|\,(e_{\text{chosen}},e_{\text{rejected}})\in\{(e_{\text{exp}},e_{\text{tgt}}),(e_{\text{exp}},e_{\text{fail}})\}\ \cup\ (\mathcal{F}_{\text{tgt}}\times\mathcal{F}_{\text{fail}})\Big\},$ (8)
where $e_{\text{exp}}$ denotes expert trajectories, $e_{\text{tgt}}$ target-predicted trajectories, and $e_{\text{fail}}$ failure-agent trajectories. Here, $\mathcal{F}_{\text{tgt}}=\{e_{\text{tgt}}\mid r(u,e_{\text{tgt}})<1\}$ and $\mathcal{F}_{\text{fail}}=\{e_{\text{fail}}\mid r(u,e_{\text{fail}})<1\}$ denote the sets of failed trajectories generated by the target and failure agents, respectively.
The target agent is optimized with a weighted DPO objective (dpo) together with an auxiliary supervised fine-tuning (SFT) loss on the chosen trajectories:
$\mathcal{L}_{\text{target}}=\lambda_{\text{DPO}}\,\mathcal{L}_{\text{DPO}}+\lambda_{\text{SFT}}\,\mathbb{E}_{(u,e_{\text{chosen}})\sim\mathcal{D}_{\text{tgt}}}\big[-\log\pi_{\theta}(e_{\text{chosen}}\mid u)\big].$ (9)
As noted by yuan, DPO alone maximizes relative preference margins but can become unstable, since the space of chosen trajectories is much smaller than that of rejected ones. This imbalance may lead the model to over-penalize rejected samples while insufficiently reinforcing preferred ones. To stabilize training, we introduce the auxiliary SFT loss on the chosen trajectories, which grounds the policy toward high-reward behaviors. For all datasets, we set $\lambda_{\text{DPO}}=1.0$ and $\lambda_{\text{SFT}}=0.1$ as the default weighting.
Additionally, since expert–prediction pairs may arise from both the target agent’s and the failure agent’s trajectories, we assign per-pair weights such that the total weight of all expert–prediction pairs generated for a given instruction sums to $1.0$. This prevents these pairs from dominating the training signal while still encouraging alignment toward high-reward behaviors. For failure–failure pairs, we apply DPO exclusively by setting $\lambda_{\text{SFT}}=0$, relying purely on preference optimization to avoid introducing misleading supervision from incorrect trajectories.
This design prevents over-counting of expert signals while ensuring that failure pairs receive full emphasis. Failure–failure comparisons are especially valuable, as they capture subtle distinctions between suboptimal behaviors, thereby sharpening the agent’s decision boundaries. As training alternates with the failure agent, the two agents form an implicit arms race: the failure agent generates increasingly challenging negatives, while the target agent learns to overcome them. This co-evolutionary loop not only enhances robustness within the training domain but also improves generalization to unseen environments where expert supervision is limited.
@@ 5.1 Experimental settings
We conduct experiments on three representative benchmarks: WebShop for web navigation, ScienceWorld for scientific reasoning, and InterCodeSQL for interactive SQL querying. All three environments provide continuous final rewards in $[0,1]$, enabling fine-grained evaluation of task completion. Expert trajectories are collected through a combination of human annotations and GPT-4–assisted generation in the ReAct format (react), with additional filtering based on final rewards to ensure quality. We present the overview statistics in Table 1. Also, Example trajectory samples for each dataset and further details of the environments and trajectory collection process are provided in Appendix A.
@@
We adopt Llama-2-7B-Chat (llama2) and Qwen3-4B-Instruct-2507 (qwen3). All models are optimized with AdamW (adamw), and each training phase is performed for 3 epochs with co-evolution iterations set to 3 for WebShop and ScienceWorld and 5 for InterCodeSQL. All other hyperparameters are kept identical across datasets to ensure fair comparison. Experiments are conducted on 8 NVIDIA H100 GPUs with 80GB memory, and further implementation details are provided in Appendix B.
@@
We compare our framework with standard imitation learning and several strong post-imitation baselines following eto. Supervised fine-tuning (SFT) (sft1; sft2) trains agents via behavioral cloning on expert trajectories and serves as the base policy for other methods. Rejection Fine-Tuning (RFT) (rft) augments the expert dataset with success trajectories identified by rejection sampling, while Proximal Policy Optimization (PPO) (ppo) directly optimizes the SFT policy with reinforcement learning to maximize task rewards. We additionally include ETO (eto), which applies DPO over expert-versus-predicted pairs without the co-evolution mechanism, thus serving as a DPO-only baseline for isolating the contribution of failure-agent learning. For reference, we also report results from GPT-3.5-Turbo (gpt3.5) and GPT-4 (gpt4) with in-context learning. We report average reward as the primary evaluation metric.
Figure 2: Diversity analysis
Figure 3: Ablation on $\lambda_{\text{SFT}}$.
@@ 5.2.1 Quantitative Analysis
The distribution of generated trajectories
We separate negative and hard-negative trajectories using a reward threshold of 0.6. Although a higher threshold would be preferable for identifying ideal hard negatives, trajectories with reward above 0.7–0.8 appear in fewer than 1% of cases in current self-improving agents. The resulting statistics are summarized in Table 2. Across all benchmarks, our method generates substantially more negative and hard-negative trajectories than ETO. Specifically, negative trajectories increase by about 9.5% in WebShop, 16.7% in ScienceWorld, and 9.0% in InterCodeSQL. Hard negatives similarly increase by 2.3%, 8.7%, and 4.3%, respectively. These consistent gains indicate that the failure agent effectively expands the informative failure space.
@@
We measure the diversity of failure trajectories by embedding each generated trajectory using the LLM2Vec–Meta-Llama-3-8B-Instruct-mntp llm2vec as a sentence-embedding model and computing the average pairwise distance among trajectories generated per instruction during training. As shown in Figure 2, the failure agent produces a larger number of trajectories with higher distances compared to ETO, and the boxplot demonstrates a notably larger mean and standard deviation. This confirms that the failure agent explores a broader and more diverse failure space rather than overfitting to specific modes.
Additionally, although the failure agent is trained only on failure trajectories, we observe that it practically does not collapse in our experiments. Its joint training on the target agent’s failure trajectories, which often contain partial progress toward the goal, allows it to maintain a comparable level of competence and prevents degeneration into trivial failures.
5.2.2 Qualitative Analysis
We qualitatively examine the role of hard negatives through representative examples, focusing on (i) the quality of the hard negatives and (ii) the effect of hard negatives.
The quality of the hard negatives
We analyze the WebShop task of buying a machine-washable curtain. As shown in the example below, ETO produces a shallow failure that ignores key constraints such as washability and price. In contrast, our method generates a more structured failure: the agent systematically filters mismatching candidates, checks all relevant attributes, verifies constraints, and arrives at a nearly correct choice. Although still unsuccessful, the trajectory exhibits a coherent decision process and serves as a prototypical hard negative, reflecting substantially higher quality than the failures produced by ETO. Additional examples are provided in Appendix C.
The effect of hard negatives
We examine how these trajectories influence the target agent’s behavior. As illustrated in the ScienceWorld example of growing a lemon shown below, ETO fails to make meaningful progress: after planting seeds, it repeatedly issues waiting or invalid actions and does not attempt soil preparation, tool use, or environment control. These shallow failures provide limited supervision for preference optimization.
In contrast, our method produces hard-negative failures that attempt multiple sub-skills and carry out most of the required pipeline, including navigation, soil collection, soil preparation, planting, and greenhouse regulation. Although these trajectories still fail to achieve the final goal, they exhibit coherent multi-step behavior that provides substantially richer supervision for DPO. As a result, the target agent acquires these sub-skills more effectively and completes the task more reliably in subsequent iterations. A full version of this example is provided in Appendix C.
5.3 Main Results
We evaluate our framework on three challenging multi-step decision-making benchmarks: WebShop for web navigation, ScienceWorld for scientific reasoning, and InterCodeSQL for interactive SQL querying. All benchmarks use a normalized reward in $[0,1]$, enabling a fine-grained comparison of task success.
Results on Llama-2.
Table 3 summarizes the results on Llama-2 across WebShop, ScienceWorld, and InterCodeSQL. Under in-context learning without fine-tuning, GPT-4 and GPT-3.5-Turbo achieve $63.2$ and $62.4$ on WebShop, but their performance drops sharply on ScienceWorld and InterCodeSQL, illustrating the limitations of prompt-only adaptation for reasoning-intensive or previously unseen domains.
Fine-tuning methods substantially improve performance by aligning models with environment-specific interaction patterns. Among these methods, ETO serves as a strong overall baseline with an average reward of $58.3$ across benchmarks. Please note that the results of PPO and RFT on ScienceWorld are taken from eto; IPR.
Our method achieves an average reward of $64.1$, outperforming ETO ($58.3$) by $5.8\%$. The improvements are consistent across all benchmarks: WebShop ($3.7\%$), seen ScienceWorld ($4.1\%$), unseen ScienceWorld ($6.5\%$), and InterCodeSQL ($4.4\%$). The largest gain appears on the unseen ScienceWorld split, indicating that leveraging diverse, near-success hard negatives leads to substantially stronger generalization to unfamiliar scientific environments.
Results on Qwen3.
Table 4 reports results using Qwen3. Compared to Llama-2, Qwen3 shows a stronger SFT baseline on WebShop ($63.9$) but exhibits slightly limited performance on ScienceWorld and particularly low performance on InterCodeSQL. This reflects differences in the pretrained models’ prior knowledge, including weaker grounding in scientific reasoning and minimal prior exposure to SQL-style interactive querying.
Fine-tuning with ETO provides substantial gains, achieving an average reward of $59.5$ across benchmarks. Our method further improves performance to $66.3$, surpassing ETO by $6.8\%$. The improvements are consistent across all tasks: WebShop ($6.8\%$), seen ScienceWorld ($6.5\%$), unseen ScienceWorld ($3.3\%$), and InterCodeSQL ($10.3\%$). Notably, the gains on both seen and unseen ScienceWorld splits indicate that the proposed failure-agent framework generalizes well even with smaller model scales.
Overall, these results demonstrate that our approach provides consistent and substantial improvements over ETO across architectures, highlighting the robustness of leveraging diverse, near-success hard negatives produced through co-evolution.
Comparison with DART-style Multi-Sampling.
Recent preference-optimization methods apply multi-sample rollouts to estimate instruction-level difficulty and select informative trajectories (e.g.,DART-style data filtering). To strengthen the RFT baseline, we incorporate this approach by using a multi-sample budget of $N=10$ and computing instruction-level difficulty with $n=5$ rollouts. Concretely, for each instruction we first generate $n=5$ trajectories and estimate a difficulty score based on the fraction of successful rollouts. This difficulty score is then multiplied by the multi-sample budget ($N=10$) to determine how many additional rollouts to generate for each instruction. The resulting successful trajectories are collected and used for RFT-style SFT training.
However, the effect of DART remains limited: RFT achieves 60.89 on WebShop, and RFT+DART increases this only to 61.90 on Llama-2. This limited gain reflects the nature of our target tasks. Multi-sampling typically excels in domains such as mathematics, where pretrained LLMs hold substantial prior knowledge and multiple rollouts are likely to contain correct trajectories. In contrast, our benchmarks require grounded, environment-specific reasoning in web navigation, scientific procedures, and interactive SQL, where pretrained models have weaker priors. As a result, multi-sampled rollouts rarely produce successful trajectories, substantially reducing the effectiveness of difficulty-based selection in this setting.
5.4 Ablation Study
Ablation on Varying $\lambda_{\text{SFT}}$.
We conduct an ablation study to examine the sensitivity of our method to the SFT weight $\lambda_{\text{SFT}}$. As shown in Figure 3, the average reward remains stable across a wide range of values (0.01, 0.1, 0.5), suggesting that the proposed method is reasonably robust as long as the SFT term is kept weaker than DPO.
Parameter-Matched Ablation on the Failure Agent.
@@ 5.4 Ablation Study
Table 3: Main results.
Adaptation Models WebShop ScienceWorld InterCodeSQL Avg.
@@
Fine-tuning Qwen3-4B-Instruct-2507 + SFT 63.9 43.6 40.8 12.7 40.3
@@
Table 4: Results on Qwen3-4B-Instruct-2507
@@ 6 Conclusion
We introduced a co-evolving agents framework where a target agent and a failure agent learn in alternating phases and improve through mutual interaction. By training the failure agent on failure–failure preferences, it generates near-success failures that serve as informative hard negatives. Incorporating these trajectories into the target agent’s preference optimization sharpens decision boundaries and improves robustness and generalization. Experiments across WebShop, ScienceWorld, and InterCodeSQL demonstrate consistent gains in diverse domains, underscoring that failures, when refined into structured signals, can be transformed into valuable resources for self-improving agents. We hope our findings facilitate more principled handling of failure trajectories, ultimately contributing to the advancement of the next generation of self-improving agents.

### v4

@@ 4.3 Co-evolutionary Training
In our framework, the target agent and the failure agent improve jointly through alternating optimization. The failure agent provides increasingly informative hard negatives, while the target agent leverages them to refine its preference landscape. This co-evolving process creates a feedback loop that strengthens both agents and drives improved generalization.
Preference Dataset.
The target agent is updated using a preference dataset that combines expert demonstrations with failures from both agents. Specifically, we construct $\mathcal{D}_{\text{tgt}}$ using (i) expert–target comparisons, (ii) expert–failure-agent comparisons, and (iii) failure–failure comparisons between the two agents. Formally,
$\mathcal{D}_{\text{tgt}}={}$ $\{(u,e_{\text{exp}},e^{-})\mid e^{-}\in\{e_{\text{tgt}},e_{\text{fail}}\}\}$
$\cup\;\{(u,e^{+},e^{-})\mid(e^{+},e^{-})\in\mathcal{F}_{\text{tgt}}\times\mathcal{F}_{\text{fail}}\},$
where $e_{\text{exp}}$ denote expert trajectories. This preference dataset provides more informative comparisons through interactions with hard negatives, enabling the target agent to learn a more coherent preference landscape.
Preference Optimization.
The target agent is optimized with a weighted DPO objective (Rafailov et al., 2023) together with an auxiliary supervised fine-tuning (SFT) loss on the chosen trajectories:
$\mathcal{L}_{\text{target}}(\theta_{t})={}$ $\lambda_{\text{DPO}}\,\mathcal{L}_{\text{DPO}}(\theta_{t})$
$\quad+\lambda_{\text{SFT}}\,\mathbb{E}_{(u,e^{+})\sim\mathcal{D}_{\text{tgt}}}\big[-\log\pi_{\theta_{t}}(e^{+}\mid u)\big].$
As noted by Yuan et al. (2025a), DPO alone maximizes relative preference margins but can become unstable, since the space of chosen trajectories is much smaller than that of rejected ones. This imbalance may lead the model to over-penalize rejected samples while insufficiently reinforcing preferred ones. We therefore include an auxiliary SFT term to anchor the policy toward high-reward behaviors and stabilize optimization. Throughout all experiments, we use fixed weights of $\lambda_{\text{DPO}}=1.0$ and $\lambda_{\text{SFT}}=0.1$.
As training alternates between the target agent and the failure agent, the two models form an implicit arms race: the failure agent continually produces harder and more informative negatives, while the target agent learns to overcome them. This mutual pressure encourages both agents to explore more fine-grained distinctions within the task space. Through this co-evolving process, the target agent acquires a more coherent preference landscape and develops improved generalization.
@@ 5.1 Experimental settings
We conduct experiments on three representative benchmarks: WebShop for web navigation, ScienceWorld for scientific reasoning, and InterCodeSQL for interactive SQL querying. All three environments provide continuous final rewards in $[0,1]$, enabling fine-grained evaluation of task completion. Expert trajectories are collected through a combination of human annotations and GPT-4–assisted generation in the ReAct format (Yao et al., 2023), with additional filtering based on final rewards to ensure quality. We present the overview statistics in Table 1.Also, Example trajectory samples for each dataset and further details of the environments and trajectory collection process are provided in Appendix A.
@@
We adopt Llama-2-7B-Chat (Touvron et al., 2023), Llama-2-13B-Chat (Touvron et al., 2023), and Qwen3-4B-Instruct-2507 (Yang et al., 2025). All models are optimized with AdamW (Loshchilov and Hutter, 2017), and each training phase is performed for 3 epochs with co-evolution iterations set to 3 for WebShop and ScienceWorld and 5 for InterCodeSQL. All other hyperparameters are kept identical across datasets to ensure fair comparison. Experiments are conducted on 8 NVIDIA H100 GPUs with 80GB memory, and further implementation details are provided in Appendix B.
@@
We compare our framework with standard imitation learning and several strong post-imitation baselines following Song et al. (2024). Supervised fine-tuning (SFT) (Chen et al., 2023; Zeng et al., 2024) trains agents via behavioral cloning on expert trajectories and serves as the base policy for other methods. Rejection Fine-Tuning (RFT) (Yuan et al., 2023) augments the expert dataset with success trajectories identified by rejection sampling, and we further strengthen this baseline using DART-style difficulty-aware multi-sampling (Tong et al., 2024), which allocates additional rollouts to instructions estimated as more difficult. Proximal Policy Optimization (PPO) (Schulman et al., 2017) directly optimizes the SFT policy with reinforcement learning to maximize task rewards. We additionally include ETO (Song et al., 2024), which applies DPO over expert-versus-predicted pairs without the co-evolution mechanism, thus serving as a DPO-only baseline for isolating the contribution of failure-agent learning. For reference, we also report results from GPT-3.5-Turbo (Ouyang et al., 2022) and GPT-4 (Achiam et al., 2023) with in-context learning. We report average reward as the primary evaluation metric.
@@ 5.2.1 Quantitative Analysis
Distribution of generated trajectories.
We separate negative and hard-negative trajectories using a reward threshold of 0.6. Higher thresholds (0.7-0.8) would be ideal but occur in fewer than 1% of cases in current self-improving agents. Table 2 summarizes the resulting statistics. Across all benchmarks, our failure agent $\text{Fail}_{\text{ours}}$ produces substantially more negative and hard-negative trajectories than ETO: negative trajectories increase by 9.5% (WebShop), 16.7% (ScienceWorld), and 9.0% (InterCodeSQL), while hard negatives increase by 2.3%, 8.7%, and 4.3%, respectively. These consistent gains show that $\text{Fail}_{\text{ours}}$ effectively expands the informative failure space.
@@
We assess trajectory diversity by embedding each generated failure using LLM2Vec–Meta-Llama-3-8B-Instruct-mntp (BehnamGhader et al., 2024) and computing the average pairwise distance among trajectories generated per instruction. As shown in Figure 2, the failure agent produces substantially more trajectories with larger pairwise distances than ETO, with the boxplot exhibiting a higher mean and variance. This indicates that the failure agent explores a broader and more diverse failure space rather than collapsing to narrow modes.
@@
Table 3: Results on Llama-2-7B-Chat.
Models WebShop ScienceWorld InterCodeSQL Avg.
@@
Qwen3-4B-Instruct-2507 + SFT 63.9 43.6 40.8 12.7 40.3
@@
Table 4: Results on Qwen3-4B-Instruct-2507.
(a) Distance distribution.
(b) Boxplot.
Figure 2: Diversity analysis.
5.2.2 Qualitative Analysis
We qualitatively examine the role of hard negatives through representative examples, focusing on (i) their quality and (ii) their effect on learning.
The quality of the hard negatives
We analyze the WebShop task of purchasing a machine-washable curtain. As shown in the example, ETO produces a shallow failure that overlooks key constraints such as washability and price. In contrast, our method generates a more structured near-miss: the agent filters mismatching items, checks relevant attributes, verifies constraints, and ends with an almost correct decision. Although still a failure, the trajectory reflects a coherent decision process and exemplifies a high-quality hard negative. Additional examples are provided in Appendix D.
The effect of hard negatives
We next examine how hard negatives influence the target agent’s behavior. In the ScienceWorld task of growing a lemon, ETO produces shallow failures: after planting seeds, the agent repeatedly issues waiting or invalid actions and never attempts soil preparation, tool use, or environment control, offering little usable supervision. In contrast, our method generates hard negatives that attempt multiple sub-skills and execute most of the required pipeline including navigation, soil collection, soil preparation, planting, and greenhouse regulation. Although still unsuccessful, these trajectories provide coherent multi-step behavior and much richer preference signals. As a result, the target agent acquires the underlying sub-skills more effectively and succeeds more reliably in later iterations. A full example is provided in Appendix D.
5.3 Main Results
We evaluate our framework on three challenging multi-step decision-making benchmarks: WebShop for web navigation, ScienceWorld for scientific reasoning, and InterCodeSQL for interactive SQL querying. All benchmarks use a normalized reward in $[0,1]$, allowing fine-grained comparison of task performance.
Results on Llama-2-7B.
Table 3 reports results across the three benchmarks. Under in-context learning, GPT-4 and GPT-3.5-Turbo perform reasonably on WebShop but degrade substantially on ScienceWorld and InterCodeSQL, highlighting the limitations of prompt-only adaptation in grounded or reasoning-heavy environments. Fine-tuning methods improve performance by aligning models with environment-specific interaction patterns. Among them, ETO is a strong baseline with an average reward of $58.3$ across tasks (PPO and RFT results on ScienceWorld are taken from Song et al. (2024); Xiong et al. (2024)).
Our method achieves an average reward of $64.1$, outperforming ETO by $5.8\%$. Gains are consistent across all benchmarks: WebShop ($3.7\%$), seen ScienceWorld ($4.1\%$), unseen ScienceWorld ($6.5\%$), and InterCodeSQL ($4.4\%$). The largest improvement on unseen ScienceWorld suggests that diverse near-success hard negatives significantly enhance generalization to unfamiliar scientific environments.
Results on Llama-2-13B.
Table 5 shows that our method also improves performance on the larger Llama-2-13B model. Across WebShop and the ScienceWorld splits, our approach consistently outperforms both SFT and ETO, demonstrating that the benefits of co-evolutionary training scale beyond the 7B setting.
Models WebShop ScienceWorld
Seen Unseen
Llama-2-13B-chat + SFT 59.3 64.5 56.1
Llama-2-13B-chat + ETO 65.0 72.6 65.3
Llama-2-13B-chat + Ours 69.2 74.5 65.5
Table 5: Results on Llama-2-13B-chat.
Results on Qwen3.
Table 4 presents the results on Qwen3. Although Qwen3 provides a strong SFT baseline on WebShop, its performance on ScienceWorld and especially InterCodeSQL is relatively limited, reflecting weaker grounding in scientific reasoning and interactive SQL.
Fine-tuning with ETO substantially boosts performance, and our method provides further consistent gains across all tasks: approximately $6.8\%$ on WebShop, $6.5\%$ on seen ScienceWorld, $3.3\%$ on unseen ScienceWorld, and $10.3\%$ on InterCodeSQL. These improvements demonstrate that our failure-agent framework yields substantial and reliable gains across architectures, highlighting the effectiveness of leveraging diverse, near-success hard negatives.
Comparison with DART-style Multi-Sampling.
We also strengthen the RFT baseline by incorporating DART-style multi-sampling, using an initial $n=5$ rollouts to estimate instruction-level difficulty and allocating a total budget of $N=10$ samples per instruction. Successful trajectories from this process are added to the RFT training set.
However, the effect of DART is modest in our setting: on WebShop, RFT improves from $60.89$ to only $61.90$ with DART. This limited impact stems from the nature of our benchmarks. Unlike domains such as mathematics—where pretrained LLMs already possess strong priors and multi-sampling frequently yields correct trajectories—our tasks require grounded, environment-specific reasoning in web navigation, scientific procedures, and interactive SQL. Consequently, successful rollouts are rare, reducing the usefulness of difficulty-based selection.
5.4 Ablation Study
Ablation on Varying $\lambda_{\text{SFT}}$.
We examine the sensitivity of our method to the SFT weight $\lambda_{\text{SFT}}$ by evaluating three settings ($0.01$, $0.1$, $0.5$) on the WebShop benchmark using Llama-2-7B-chat. As shown in Table 6, the resulting average rewards (66.98, 66.70, 66.50) vary only slightly, indicating that the method is robust as long as the SFT term remains weaker than DPO.
$\lambda_{\text{SFT}}{=}0.01$ $\lambda_{\text{SFT}}{=}0.1$ $\lambda_{\text{SFT}}{=}0.5$
Avg. Reward 67.98 66.70 66.50
Table 6: Ablation on the SFT weight $\lambda_{\text{SFT}}$.
Parameter-Matched Ablation on the Failure Agent.
@@ 6 Conclusion
We presented a co-evolving agents framework in which a target agent and a failure agent improve jointly through alternating preference-optimization phases. The failure agent, trained solely on failure–failure comparisons, learns a fine-grained failure landscape and autonomously produces near-success hard negatives. Incorporating these trajectories into the target agent’s preference optimization provides sharper and more informative contrasts than standard expert–failure supervision, leading to a more coherent preference landscape. Experiments on WebShop, ScienceWorld, and InterCodeSQL demonstrate consistent improvements across domains and model architectures, showing that refining failures into structured supervision is an effective mechanism for strengthening generalization in self-improving agents. We hope these findings encourage more principled use of failure trajectories as a first-class training signal in the development of next-generation self-improving agents.
Impact Statement
This work aims to advance the development of more robust self-improving agents by leveraging failure trajectories as learning signals. While strengthening autonomous decision-making systems can have broad downstream implications, including potential misuse or reinforcement of existing biases, our method does not introduce new data modalities or rely on sensitive information. As with any system that adapts autonomously, appropriate safeguards and careful evaluation are recommended in safety-critical applications. Beyond these general considerations, we do not identify additional societal impacts requiring specific attention.